# Prompting Depth Anything for 4K Resolution Accurate Metric Depth Estimation

### v2

@@ Abstract
Prompts play a critical role in unleashing the power of language and vision foundation models for specific tasks. For the first time, we introduce prompting into depth foundation models, creating a new paradigm for metric depth estimation termed Prompt Depth Anything. Specifically, we use a low-cost LiDAR as the prompt to guide the Depth Anything model for accurate metric depth output, achieving up to 4K resolution. Our approach centers on a concise prompt fusion design that integrates the LiDAR at multiple scales within the depth decoder. To address training challenges posed by limited datasets containing both LiDAR depth and precise GT depth, we propose a scalable data pipeline that includes synthetic data LiDAR simulation and real data pseudo GT depth generation. Our approach sets new state-of-the-arts on the ARKitScenes and ScanNet++ datasets and benefits downstream applications, including 3D reconstruction and generalized robotic grasping.
Figure 1: Illustration and capabilities of Prompt Depth Anything. (a) Prompt Depth Anything is a new paradigm for metric depth estimation, which is formulated as prompting a depth foundation model with a metric prompt, specifically utilizing a low-cost LiDAR as the prompt. (b) Our method enables consistent depth estimation, addressing the limitations of Metric3D v2 [26] that suffer from inaccurate scale and inconsistency. (c) It achieves accurate 4K accurate depth estimation, significantly surpassing ARKit LiDAR Depth (240 $\times$ 320).
Corresponding author: Sida Peng
1 Introduction
High-quality depth perception is a fundamental challenge in computer vision and robotics. Recent monocular depth estimation has experienced a significant leap by scaling the model or data, leading to the flourishing of depth foundation models [75, 76, 30, 19]. These models demonstrate strong abilities in producing high-quality relative depth, but suffer from scale ambiguity, hindering their practical applications in autonomous driving and robotic manipulation, etc. Therefore, significant efforts have been made to achieve metric depth estimation, by either finetuning depth foundation models [6, 20] on metric datasets or training metric depth models with image intrinsics as additional inputs [46, 80, 8, 26]. However, neither of them can address the problem properly, as illustrated in Fig. 1(b).
@@ 1 Introduction
  • A concise prompt fusion architecture for depth foundation models, a scalable data pipeline, and an edge-aware depth loss to train Prompt Depth Anything.
  • State-of-the-art performance on depth estimation benchmarks [78, 3], showing the extensibility of replacing depth foundation models and LiDAR sensors, and highlighting benefits for several downstream applications including 3D reconstruction and robotic object grasping.
2 Related Work
@@
Traditional methods [54, 25] rely on hand-crafted features for depth estimation. With the advent of deep learning, this field has seen significant advancements. Early learning-based approaches [16, 15] are often limited to a single dataset, lacking generalization capabilities. To enhance generalization, diverse datasets [34, 77, 13, 66, 63, 62, 61, 65], affine-invariant loss [47], and more powerful network architectures [48] have been introduced. More recently, latent diffusion models [52], pre-trained on extensive image generation tasks, have been applied to depth estimation [30, 22]. These models exhibit good generalization, estimating relative depth effectively, though they remain scale-agnostic. To achieve metric depth estimation, early methods either model the problem as global distribution classification [18, 4, 5, 37] or fine-tune a depth model on metric depth datasets [6, 35, 36]. Recent methods [80, 20, 79, 26, 46] discuss the ambiguity in monocular metric depth estimation and address it by incorporating camera intrinsic parameters. Although recent methods [80, 26, 46, 8, 76, 30, 22] exhibit strong generalization ability and claim to be depth foundation models [8, 76, 26, 19], metric depth estimation remains a challenge as shown in Fig. 1(b). We seek to address this challenge by prompting the depth foundation models with a metric prompt, inspired by the success of prompting in vision and vision-language models [40, 39, 83].
@@
Obtaining dense depth information through active sensors typically demands high power consumption [68, 70, 69, 71, 72, 10]. A more practical approach involves utilizing a low-power active sensor to capture sparse depth, which can then be completed into dense maps. Many studies investigate methods to fill in sparse depth data. Early works rely on filter-based [23, 28, 32] and optimization-based [17, 74] techniques for depth completion. More recent studies [59, 84, 12, 11, 58, 41, 57, 72, 38, 14, 21] adopt learning-based approaches for depth completion. Typically, these methods are not tested on real indoor LiDAR data but rather on simulated sparse lidar for depth datasets such as NYUv2 [16] to reconstruct complete depth. This is because real testing setups require both low-power and high-power LiDAR sensors. More recent works have collected both low-power and high-power LiDAR data. To collect such data, DELTA [33] builds a suite to collect data using L5 and Intel RealSense 435i, while three other datasets [3, 78, 50] are collected using iPhone LiDAR and FARO LiDAR. We focus on the latter, as iPhone is widely available. A recent work similar to ours is Depth Prompting [43]. Our approach differs in that we use a network to take sparse depth as a prompt for the depth foundation model, achieving specific output. In contrast, they fuse sparse depth with features from the depth foundation model to post-process the foundation model output, which does not constitute prompting a foundation model.
3 Method
@@
Monocular depth estimation models [75, 76, 8] are becoming depth foundation models for their generalization ability obtained from large-scale data. However, due to the inherent ambiguities, they cannot achieve high accuracy on metric depth estimation as shown in Fig. 1(b). Inspired by the success of prompting for vision [31, 52, 40] and language [1] foundation models, we propose Prompt Depth Anything prompting the depth foundation model with a metric prompt to achieve metric depth estimation. We take the low-cost LiDAR as the metric prompt in this work, as it has recently been integrated into lots of smartphones, making this setup highly practical. To be specific, we aim to prompt the depth foundation model to unleash its power for accurate metric depth estimation.
3.1 Preliminary: Depth Foundation Model
@@
Current depth foundation models [75, 76, 80, 7] generally share similar network structures of DPT [48] networks. Specifically, given an image $\mathbf{I}\in\mathbb{R}^{C\times H\times W}$, they take a vision transformer (ViT) with multiple stages to extract tokenized image features $\{\mathbf{T}_{i}\}$, where $\mathbf{T}_{i}\in\mathbb{R}^{C_{i}\times(\frac{H}{p}\times\frac{W}{p}+1)}$ represents the feature map at stage $S_{i}$, $D_{i}$ is the feature dimension at stage $S_{i}$, and $p$ is the patch size. The DPT decoder reassembles features from different stages into image-like representations $\mathbf{F}_{i}\in\mathbb{R}^{D_{i}\times\frac{H}{p}\times\frac{W}{p}}$ with the reassemble operation [48]. Finally, a sequence of convolutional blending steps are applied to merge features ${\mathbf{F}_{i}}$ across different stages, predicting a dense depth map $\mathbf{D}\in\mathbb{R}^{H\times W}$.
We note that there exists another line of depth foundation models [30, 19, 22] that use the image diffusion model [53] to estimate depth maps. Due to the high computational cost of diffusion models, we only consider DPT-based depth foundation models [76, 8] as our base model for real-time performance in this work.
3.2 Prompt Depth Anything
@@
In this section, we seek to find a concise way to incorporate a low-cost LiDAR (i.e., a low-resolution and noisy depth map) as a prompt into the depth foundation model. To this end, we propose a concise prompt fusion architecture tailored for the DPT-based [48] depth foundation models to integrate low-resolution depth information. As shown in Fig. 2(a), the prompt fusion architecture integrates low-resolution depth information at multiple scales within the DPT Decoder. Specifically, for each scale $S_{i}$ in the DPT Decoder, a low-resolution depth map $\mathbf{L}\in\mathbb{R}^{1\times H_{\mathbf{L}}\times W_{\mathbf{L}}}$ is firstly bilinearly resized to match the spatial dimensions of the current scale $\mathbb{R}^{1\times H_{i}\times W_{i}}$. Then, the resized depth map is passed through a shallow convolutional network to extract depth features. After that, the extracted features are projected to the same dimension as the image features $\mathbf{F}_{i}\in\mathbb{R}^{C_{i}\times H_{i}\times W_{i}}$ using a zero-initialized convolutional layer. Finally, the depth features are added to the DPT intermediate features for depth decoding. The illustration of this block design is shown in Fig. 2.
@@
The proposed design has the following advantages. Firstly, it introduces only 5.7% additional computational overhead (1.789 TFLOPs v.s. 1.691 TFLOPs for a $756\times 1008$ image) to the original depth foundation model, and effectively addresses the ambiguity issue inherent in the depth foundation model as demonstrated in Tab. 3(b). Secondly, it fully inherits the capabilities of the depth foundation model because its encoder and decoder are initialized from the foundation model [76], and the proposed fusion architecture is zero-initialized, ensuring that the initial output is identical to that of the foundation model. We experimentally verify the importance of inheriting from a pretrained depth foundation model as shown in Tab. 3(c).
@@
Inspired by conditional image generation methods [83, 44, 29], we also explore various potential prompt conditioning designs into the depth foundation model. Specifically, we experimented with the following designs: a) Adaptive LayerNorm [45, 29] which adapts the layer normalization parameters of the encoder blocks based on the conditioning input, b) CrossAttention [60] which injects a cross attention block after each self-attention block and integrates the conditioning input through cross-attention mechanisms, and c) ControlNet [83] which copies the encoder blocks and inputs control signals to the copied blocks to control the output depth. As shown in Tab. 3(d,e,f), our experiments reveal that these designs do not perform as well as the proposed fusion block. A plausible reason is that they are designed to integrate cross-modal information (e.g., text prompts), which does not effectively utilize the pixel alignment characteristics between the input low-res LiDAR and the output depth. We detail these optional designs in the supp.
@@
3.3 Training Prompt Depth Anything
@@
A LiDAR depth map is low-resolution and noisy. The naive approach for simulating it is to directly downsample the synthetic data depth map. However, this method leads to the model learning depth super-resolution, as shown in Fig. 3, meaning that the model does not correct the LiDAR noise. To simulate the noise, we introduce a sparse anchor interpolation method. Specifically, we first downsample the GT depth map to low-resolution ($192\times 256$, exactly the depth resolution of iPhone ARKit Depth). Then we sample points on this depth map using a distorted grid with a stride (7 in practice). The remaining depth values are interpolated from these points using RGB similarity with KNN. As shown in Fig. 3, it effectively simulates LiDAR noise and results in better depth prediction. We provide visualization results of the simulated LiDAR in the supp.
@@
We also add real data [78] to our training data. The annotated depth in ScanNet++ [78] is re-rendered from a mesh scanned by a high-power LiDAR sensor (FARO Focus Premium laser scanner). Due to the presence of many occlusions in the scene, several scan positions (typically 4 in a medium-sized scene in ScanNet++) result in an incomplete scanned mesh, leading to depth maps with numerous holes and poor edge quality, as illustrated in Fig. 2(b). Motivated by the success of reconstruction methods [2, 42], we propose using Zip-NeRF [2] to recover high-quality depth maps. Specifically, we train Zip-NeRF for each scene in ScanNet++ and re-rendered pseudo GT depth. To provide Zip-NeRF with high-quality and dense observations, we detect unblurred frames in Scannet++iPhone videos, and additionally utilize DSLR videos to provide high-quality dense-view images.
@@
Although Zip-NeRF can generate high-quality edge depth, reconstructing textureless and reflective regions remains challenging as shown in Fig. 2(b). In contrast, these areas (e.g., walls, floors, and ceilings etc.) are usually planar with few occlusions, and the annotations depth in FARO rendered depth is good in these regions. This motivates us to leverage the strengths of both. We propose an edge-aware depth loss to meet these requirements. Specifically, we use the FARO scanned mesh depth and the gradient of the pseudo GT depth to supervise output depth and the gradient of the output depth, respectively:
$\mathcal{L}_{\text{edge}}=L_{1}(\mathbf{D}_{\text{gt}},\hat{\mathbf{D}})+\lambda\cdot\mathcal{L}_{\text{grad}}(\mathbf{D}_{\text{pseudo}},\hat{\mathbf{D}}),$ (1)
@@
In practice, we set $\lambda=0.5$. The depth gradient is mainly prominent at the edges, which is exactly where the pseudo GT depth excels. The gradient loss encourages the model to learn the accurate edges from the pseudo GT depth, while the L1 loss encourages the model to learn the overall depth, ultimately leading to excellent depth prediction. We experimentally verify the effectiveness of the edge-aware depth loss in Tab. 3(j) and Fig. 3.
3.4 Implementation Details

### v3

# PromptDepthAnything++: Accurate 4K Metric Depth Estimation via Pattern-Agnostic Prompting

@@ Abstract
Prompts play a critical role in unleashing the power of language and vision foundation models for specific tasks. For the first time, we introduce prompting into depth foundation models, creating a new paradigm for metric depth estimation termed Prompt Depth Anything. Specifically, we use a low-cost LiDAR as the prompt to guide the Depth Anything model for accurate metric depth output, achieving up to 4K resolution. Our approach centers on a concise prompt fusion design that integrates the LiDAR at multiple scales within the depth decoder. To address training challenges posed by limited datasets containing both LiDAR depth and precise GT depth, we propose a scalable data pipeline that includes synthetic data LiDAR simulation and real data pseudo GT depth generation. To further extend our method to work with any prompt depth points, we propose a new prompting mechanism, which serializes the input depth points into tokens and uses self-attention to enhance image tokens from depth foundation models. Our approach sets new state-of-the-arts on 8 zero-shot depth benchmarks and benefits downstream applications, including 3D reconstruction and generalized robotic grasping. The code is available at https://github.com/DepthAnything/PromptDA.
Figure 1: Illustration and capabilities of Prompt Depth Anything. (a) Prompt Depth Anything is a new paradigm for metric depth estimation, which is formulated as prompting a depth foundation model with a metric prompt, specifically utilizing a low-cost LiDAR as the prompt. (b) Our method enables consistent depth estimation, addressing the limitations of Metric3D v2 that suffer from inaccurate scale and inconsistency. (c) It achieves accurate 4K accurate depth estimation, significantly surpassing ARKit LiDAR Depth (240 $\times$ 320).
I Introduction
High-quality depth perception is a fundamental challenge in computer vision and robotics. Recent monocular depth estimation has experienced a significant leap by scaling the model or data, leading to the flourishing of depth foundation models [92, 93, 37, 21]. These models demonstrate strong abilities in producing high-quality relative depth, but suffer from scale ambiguity, hindering their practical applications in autonomous driving and robotic manipulation, etc. Therefore, significant efforts have been made to achieve metric depth estimation, by either finetuning depth foundation models [6, 24] on metric datasets or training metric depth models with image intrinsics as additional inputs [56, 96, 8, 31]. However, neither of them can address the problem properly, as illustrated in Fig.˜1(b).
@@ I Introduction
A preliminary version of this work [46] was presented at CVPR 2025. In this extension, we first introduce a new prompting mechanism that enables foundation models to work with pattern-agnostic prompt depth points. Additionally, we include experiments on six more zero-shot depth benchmarks involving indoor [65, 16, 72], outdoor [64] and autonomous driving scenes [67, 22], further validating the effectiveness and robustness of our framework. We also compare our approach against more recent baselines involving depth completion methods [70, 74, 103, 78]. Furthermore, we investigate the potential of our prompting mechanism for multi-view stereo (MVS) reconstruction by using Structure-from-Motion (SFM) points as prompts. Surprisingly, our method outperforms leading MVS methods [25, 86] by a large margin without any further fine-tuning, demonstrating the general effectiveness of our pattern-agnostic prompting approach. The code has been made publicly available at https://github.com/DepthAnything/PromptDA, which has garnered over 800 stars.
@@
  • A concise prompt fusion architecture for prompting validation and a new pattern-agnostic prompting mechanism, a scalable data pipeline, and an edge-aware depth loss to train Prompt Depth Anything.
  • State-of-the-art depth estimation performance on benchmarks for depth upsampling [95, 3] and depth completion [65, 16, 72, 64, 22, 67], showing the extensibility of replacing different depth foundation models and LiDAR sensors.
  • State-of-the-art performance on MVS reconstruction by simply taking SfM points as prompting input without any further fine-tuning, showing the general effectiveness of our pattern-agnostic prompting approach.
  • Highlighting benefits for several downstream applications including 3D reconstruction and robotic object grasping.
II Related Work
@@
Traditional methods [63, 30] rely on hand-crafted features for depth estimation. With the advent of deep learning, this field has seen significant advancements. Early learning-based approaches [18, 17] are often limited to a single dataset, lacking generalization capabilities. To enhance generalization, diverse datasets [41, 94, 14, 81, 77, 76, 75, 80], affine-invariant loss [58], and more powerful network architectures [57] have been introduced. More recently, latent diffusion models [61], pre-trained on extensive image generation tasks, have been applied to depth estimation [37, 27, 26]. These models exhibit good generalization, estimating relative depth effectively, though they remain scale-agnostic. Pixel-Perfect Depth [85, 84] further performs diffusion generation directly in the pixel space, avoiding VAE-induced artifacts and producing flying-pixel-free depth maps. To achieve metric depth estimation, early methods either model the problem as global distribution classification [20, 4, 5, 44] or fine-tune a depth model on metric depth datasets [6, 42, 43]. Recent methods [96, 24, 97, 31, 56] discuss the ambiguity in monocular metric depth estimation and address it by incorporating camera intrinsic parameters. More recently, Depth Anything 3 [45] unifies monocular and multi-view depth estimation with a single plain transformer and a unified depth-ray representation. InfiniDepth [98] enables arbitrary-resolution and fine-grained depth estimation via neural implicit fields. SHARP [51] further leverages sharp metric depth from Depth Pro [8] for real-time monocular view synthesis. Although recent methods [96, 31, 56, 8, 93, 37, 27, 45, 26, 85, 98] exhibit strong generalization ability and claim to be depth foundation models [8, 93, 31, 21, 45], metric depth estimation remains a challenge as shown in Fig.˜1(b). We seek to address this challenge by prompting the depth foundation models with a metric prompt, inspired by the success of prompting in vision and vision-language models [49, 48, 100].
@@
Obtaining dense depth information through active sensors typically demands high power consumption [83, 88, 86, 87, 89, 11]. A more practical approach involves utilizing a low-power active sensor to capture sparse depth, which can then be completed into dense maps. Many studies investigate methods to fill in sparse depth data. Early works rely on filter-based [28, 33, 39] and optimization-based [19, 91] techniques for depth completion. More recent studies [70, 101, 12, 13, 71, 50, 68, 89, 47, 15, 25, 69] adopt learning-based approaches for depth completion. Among them, LingBot-Depth [69] employs masked depth modeling with vision transformers to refine sparse and noisy depth sensor data for robotics applications. Typically, these methods are not tested on real indoor LiDAR data but rather on simulated sparse lidar for depth datasets such as NYUv2 [18] to reconstruct complete depth. This is because real testing setups require both low-power and high-power LiDAR sensors. More recent works have collected both low-power and high-power LiDAR data. To collect such data, DELTA [40] builds a suite to collect data using L5 and Intel RealSense 435i, while three other datasets [3, 95, 59] are collected using iPhone LiDAR and FARO LiDAR. We focus on the latter, as iPhone is widely available. A recent work similar to ours is Depth Prompting [53]. Our approach differs in that we use a network to take sparse depth as a prompt for the depth foundation model, achieving specific output. In contrast, they fuse sparse depth with features from the depth foundation model to post-process the foundation model output, which does not constitute prompting a foundation model.
III Method
@@
Monocular depth estimation models [92, 93, 8] are becoming depth foundation models from large-scale data training. However, due to the inherent ambiguities, they cannot achieve high accuracy on metric depth estimation as shown in Fig.˜1(b). Inspired by the success of prompting for vision [38, 61, 49] and language [1] foundation models, we propose Prompt Depth Anything prompting the depth foundation model with a metric prompt to achieve metric depth estimation. We take the low-cost LiDAR as the metric prompt in this work, as it has recently been integrated into lots of smartphones, making this setup highly practical. To be specific, we aim to prompt the depth foundation model to unleash its power for accurate metric depth estimation.
III-A Preliminary: Depth Foundation Model
@@
Current depth foundation models [92, 93, 96, 7] generally share similar network structures of DPT [57] networks. Specifically, given an image $\mathbf{I}\in\mathbb{R}^{C\times H\times W}$, they take a vision transformer (ViT) with multiple stages to extract tokenized image features $\{\mathbf{T}_{i}\}$, where $\mathbf{T}_{i}\in\mathbb{R}^{C_{i}\times(\frac{H}{p}\times\frac{W}{p}+1)}$ represents the feature map at stage $S_{i}$, $D_{i}$ is the feature dimension at stage $S_{i}$, and $p$ is the patch size. The DPT decoder reassembles features from different stages into image-like representations $\mathbf{F}_{i}\in\mathbb{R}^{D_{i}\times\frac{H}{p}\times\frac{W}{p}}$ with the reassemble operation [57]. Finally, a sequence of convolutional blending steps are applied to merge features $\mathbf{F}_{i}$} across different stages, predicting a dense depth map $\mathbf{D}\in\mathbb{R}^{H\times W}$. We note that there exists another line of depth foundation models [37, 21, 27] that use the image diffusion model [62] to estimate depth maps. Due to the high computational cost of diffusion models, we only consider DPT-based depth foundation models [93, 8] as our base model for real-time performance in this work.
III-B Prompt Depth Anything
@@
In this section, we seek to find a concise way to incorporate a low-cost LiDAR (i.e., a low-resolution and noisy depth map) as a prompt into the depth foundation model. To this end, we propose a concise prompt fusion architecture tailored for the DPT-based [57] depth foundation models to integrate low-resolution depth information. As shown in Fig.˜2(a), the prompt fusion architecture integrates low-resolution depth information at multiple scales within the DPT Decoder. Specifically, for each scale $S_{i}$ in the DPT Decoder, a low-resolution depth map $\mathbf{L}\in\mathbb{R}^{1\times H_{\mathbf{L}}\times W_{\mathbf{L}}}$ is firstly bilinearly resized to match the spatial dimensions of the current scale $\mathbb{R}^{1\times H_{i}\times W_{i}}$. Then, the resized depth map is passed through a shallow convolutional network to extract depth features. After that, the extracted features are projected to the same dimension as the image features $\mathbf{F}_{i}\in\mathbb{R}^{C_{i}\times H_{i}\times W_{i}}$ using a zero-initialized convolutional layer. Finally, the depth features are added to the DPT intermediate features for depth decoding. The illustration of this block design is shown in Fig.˜2.
@@
The proposed design has the following advantages. Firstly, it introduces only 5.7% additional computational overhead (1.789 TFLOPs v.s. 1.691 TFLOPs for a $756\times 1008$ image) to the original depth foundation model, and effectively addresses the ambiguity issue inherent in the depth foundation model as demonstrated in Tab.˜VI(b). Secondly, it fully inherits the capabilities of the depth foundation model because its encoder and decoder are initialized from the foundation model [93], and the proposed fusion architecture is zero-initialized, ensuring that the initial output is identical to that of the foundation model. We experimentally verify the importance of inheriting from a pretrained depth foundation model as shown in Tab.˜VI(c).
@@
Inspired by conditional image generation methods [100, 54, 36], we also explore various potential prompt conditioning designs into the depth foundation model. Specifically, we experimented with the following designs: a) Adaptive LayerNorm [55, 36] which adapts the layer normalization parameters of the encoder blocks based on the conditioning input, b) CrossAttention [73] which injects a cross attention block after each self-attention block and integrates the conditioning input through cross-attention mechanisms, and c) ControlNet [100] which copies the encoder blocks and inputs control signals to the copied blocks to control the output depth. As shown in Tab.˜VI(d,e,f), our experiments reveal that these designs do not perform as well as the proposed fusion block. A plausible reason is that they are designed to integrate cross-modal information (e.g., text prompts), which does not effectively utilize the pixel alignment characteristics between the input low-res LiDAR and the output depth. We detail these optional designs in the supp.
@@
III-C Training Prompt Depth Anything
@@
A LiDAR depth map is low-resolution and noisy. The naive approach for simulating it is to directly downsample the synthetic data depth map. However, this method leads to the model learning depth super-resolution, as shown in Fig.˜3, meaning that the model does not correct the LiDAR noise. To simulate the noise, we introduce a sparse anchor interpolation method. Specifically, we first downsample the GT depth map to low-resolution ($192\times 256$, exactly the depth resolution of iPhone ARKit Depth). Then we sample points on this depth map using a distorted grid with a stride (7 in practice). The remaining depth values are interpolated from these points using RGB similarity with KNN. As shown in Fig.˜3, it effectively simulates LiDAR noise and results in better depth prediction. We provide visualization results of the simulated LiDAR in the supp.
@@
We also add real data [95] to our training data. The annotated depth in ScanNet++ [95] is re-rendered from a mesh scanned by a high-power LiDAR sensor (FARO Focus Premium laser scanner). Due to the presence of many occlusions in the scene, several scan positions (typically 4 in a medium-sized scene in ScanNet++) result in an incomplete scanned mesh, leading to depth maps with numerous holes and poor edge quality, as illustrated in Fig.˜2(b). Motivated by the success of reconstruction methods [2, 52], we propose using Zip-NeRF [2] to recover high-quality depth maps. Specifically, we train Zip-NeRF for each scene in ScanNet++ and re-rendered pseudo GT depth. To provide Zip-NeRF with high-quality and dense observations, we detect unblurred frames in Scannet++iPhone videos, and additionally utilize DSLR videos to provide high-quality dense-view images.
@@
Although Zip-NeRF can generate high-quality edge depth, reconstructing textureless and reflective regions remains challenging as shown in Fig.˜2(b). In contrast, these areas (e.g., walls, floors, and ceilings etc.) are usually planar with few occlusions, and the annotations depth in FARO rendered depth is good in these regions. This motivates us to leverage the strengths of both. We propose an edge-aware depth loss to meet these requirements. Specifically, we use the FARO scanned mesh depth and the gradient of the pseudo GT depth to supervise output depth and the gradient of the output depth, respectively:
[…]
@@
In practice, we set $\lambda=0.5$. The depth gradient is mainly prominent at the edges, which is exactly where the pseudo GT depth excels. The gradient loss encourages the model to learn the accurate edges from the pseudo GT depth, while the L1 loss encourages the model to learn the overall depth, ultimately leading to excellent depth prediction. We experimentally verify the effectiveness of the edge-aware depth loss in Tab.˜VI(j) and Fig.˜3.
III-D Pattern Agnostic Depth Prompting
Figure 4: Pipeline of pattern-agnostic depth prompting. To improve the robustness of Prompt Depth Anything against prompt depth map patterns, we propose converting prompt depth points into a sequence of tokens, thereby mitigating pattern-specific influences. We first employ an MLP to project the prompt depth points into a sequence of tokens. The resulting prompt tokens are then concatenated with the image tokens and processed through a series of self-attention layers incorporating 2D uv position encoding. After that, the image tokens are enhanced by the prompt depth points, and the enhanced image tokens are fed into a DPT Decoder to reconstruct the depth map. The depth tokens output by the self-attention layers are simply discarded.
In the previous section, we presented a systematic framework for prompting depth foundation models and practical training data construction methods. This design particularly excels at handling low-resolution prompt depth, typically for depth upsampling [3, 95], but does not address various specific prompt characteristics, such as the diverse patterns of sensors in the real world. For instance, Intel RealSense sensors exhibit different patterns: they may have random holes due to occlusion between the depth sensor and RGB sensor, or due to objects being too far away, or from scanning reflective surfaces. Without specific prompting design, previous models would produce unsatisfactory results when encountering prompt depth of extreme random patterns.
To address the challenge of diverse patterns for prompt depth in practical applications, we propose a new prompting mechanism that can accept pattern-agnostic sparse points as prompts, thereby broadening the applicability of our method. Previous methods typically design complex convolutional networks or vision transformers to extract features from sparse depth maps and then fuse this information with the image branch to output depth maps. Instead of using map-based fusion, we propose a novel fusion module that serializes the sparse depth map into a sequence of tokens for network processing. Since the network no longer needs to learn the characteristics of sparse maps (such as random holes that are unfriendly to network learning), but instead learns from serialized tokens, we can robustly handle various different point patterns.
Specifically, given a sparse depth map, we first serialize the valid depth pixels into a sequence of points $(u,v,d)$, where $u$ and $v$ are the pixel coordinates and $d$ is the depth value. Then, an MLP is used to project the depth point into a token with the same dimension as the image token. Then, the resulting depth tokens are concatenated with the image tokens extracted from the depth foundation model and passed through a series of self-attention layers. To connect the depth tokens with the image tokens, we use 2D UV encoding to encode the position of image and depth point tokens, and the encoding is added to the query and key components of the self-attention layers. After processing through the self-attention layers, the image tokens are enhanced by the depth information, while the depth tokens are discarded as they serve only as auxiliary information for enhancing the image features. The enhanced image tokens are also processed with a zero-init. MLP and then plus the original image tokens to get the final image tokens, which ensuring inheriting the foundation model ability. The final image tokens are then fed into the DPT Decoder to reconstruct the final depth map. This design enables our model to handle various sensor patterns robustly, as it learns to extract meaningful depth information from tokenized representations rather than being constrained by specific spatial patterns.
III-E Implementation Details